A cold standard air Diesel cycle has a compression ratio of 18. The heat transferred to the working fluid per cycle is 2000 kJ / kg. At the beginning of the compression process cycle, the pressure is 100 kPa and the temperature is 25 ° C. Determine (a) The pressure in each state of the cycle, (b) The cut-off ratio, (c) The thermal efficiency, (d) The net work per unit of mass and (e) the effective mean pressure.
Answer #1

Hey there Dear Respected student,I hope you're doing absolutely great, I'm too glad to answer such a nice question for you. I hope it helps you. I tried my best to provide you with a best quality answer for you . I tried to utilize my full time and knowledge for this question. Its quite easy and simple.. I believe that you will understand this in one read. Yes you can do this my dear . Please carefully go through answer till last word. And oh yes I appreciate your hunger of grabbing knowledge keep it always up like this.

So here is your answer :

Qin Wout 2 Part Win 1-2 process Y1 T2 TI TL = (11) - (1.8) 19-1 298 T2 = (290)(3-198)=946.95k ( ) = (49) (946 1.4-1 946.95 PL

n = 1-1 2011 ro 17(8-1) 14 ņ= 1-1 3.1 (1.4)(3-1-1) n=0.585 = 58.5% nee = 9, * 7 = 2000X0-585 = 11906-1 KJ ( ) 1814-1 3 VI ERT

Please Feel free to ask you doubts if any. Your suggestions are most welcomed.
Please appreciate my efforts and hardwork by upvote/thumbs up ? it will be really very helpful.


Thanks for your valuable time.
Thanks for choosing this platform. All the best for your future and career. Keep learning like this always. Stay safe stay healthy. Take a good care of yourself and family . Have a great life ahead

Thanks you so much :)